OpenStudy (anonymous):

What is power series for sin(x^2)

OpenStudy (austinl):

@zepdrix Are you familiar with a way to work to it? Without just giving the answer?

zepdrix (zepdrix):

I don't really remember power series that well. I would have to work it out the long way. :D lol I have to go to school though :( I don't have time to help right now :c

OpenStudy (austinl):

I have always just had access to tables....

OpenStudy (amistre64):

its the power series for sin, but using an x^2 instead of x

OpenStudy (amistre64):

\[sin(u)=\sum\frac{(-1)^n}{(2n+1)!}u^{2n+1}\] \[sin(x^2)=\sum\frac{(-1)^n}{(2n+1)!}x^{2(2n+1)}\]

OpenStudy (austinl):

That is what I get as well @amistre64

OpenStudy (amistre64):

was the intent to construct it as a taylor series? derivatives and such?

OpenStudy (amistre64):

spose we have a function that is a constant, plus some powers of (x-c) f = c0 + powers f' = c1 + powers' f'' = 2c2 + powers'' f''' = 3.2c3 + powers''' f'''' = 4.3.2c4 + powers'''' when the (x-c) parts are set to zero, we have a way to define the coefficients that are popping out

OpenStudy (anonymous):

yeah thanks alot, tha helpped !

OpenStudy (amistre64):

f = c0 f' = c1 f'' = 2c2 f''' = 3.2c3 f'''' = 4.3.2c4 f = c0 f' = c1 f''/2 = c2 f'''/3! = c3 f''''/4! = c4 \[c_n=f^{(n)}\frac1{n!}\]